# FinSphere: A Conversational Stock Analysis Agent Equipped with Quantitative Tools based on Real-Time Database

### v1

@@ Abstract
Current financial Large Language Models (LLMs) struggle with two critical limitations: a lack of depth in stock analysis, which impedes their ability to generate professional-grade insights, and the absence of objective evaluation metrics to assess the quality of stock analysis reports. To address these challenges, this paper introduces FinSphere, a conversational stock analysis agent, along with three major contributions: (1) Stocksis, a dataset curated by industry experts to enhance LLMs’ stock analysis capabilities, (2) AnalyScore, a systematic evaluation framework for assessing stock analysis quality, and (3) FinSphere, an AI agent that can generate high-quality stock analysis reports in response to user queries. Experiments demonstrate that FinSphere achieves superior performance compared to both general and domain-specific LLMs, as well as existing agent-based systems, even when they are enhanced with real-time data access and few-shot guidance. The integrated framework, which combines real-time data feeds, quantitative tools, and an instruction-tuned LLM, yields substantial improvements in both analytical quality and practical applicability for real-world stock analysis.
FinSphere: A Conversational Stock Analysis Agent Equipped with Quantitative Tools based on Real-Time Database
Shijie Han11footnotemark: 11,2, Changhai Zhouthanks: Equal contribution1,3, Yiqing Shen1,4, Tianning Sun1,5, Yuhua Zhou1,6, Xiaoxia Wang1, Zhixiao Yang1, Jingshu Zhang1, Hongguang Lithanks: Corresponding author: harvey2@mail.ustc.edu.cn1 1JF SmartInvest Holdings Ltd, 2Columbia University, 3Fudan University, 4Johns Hopkins University, 5University of Nottingham-Ningbo, 6Zhejiang University
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in comprehending and processing natural language, extending their influence across various domains, including finance Li et al. (2023b). By leveraging their language comprehension capabilities, these models have exhibited exceptional performance in various financial applications, including sentiment analysis Liu et al. (2024); Zhang et al. (2023a) and information extraction from unstructured financial texts Li et al. (2024); Huang et al. (2023). The advent of finance-specific LLMs such as FinBERT Yang et al. (2020); Liu et al. (2021b), BloombergGPT Wu et al. (2023), and PIXIU Xie et al. (2023) has further enhanced the capacity to process financial data effectively. These advancements have laid the foundation for developing more sophisticated financial analysis tools and shifted how investors interact with market data Krause (2023); Nie et al. (2024). These AI-powered systems have democratized access to professional-grade financial insights, enabling retail investors to obtain sophisticated financial analysis once exclusive to institutional investors.
@@
As LLM technology continues to evolve, there is an increasing expectation for these models to handle more complex financial tasks, particularly in stock analysis Yang et al. (2023b); Zhao et al. (2024). This progression has resulted in the emergence of tool-augmented agents that combine LLMs’ language capabilities with specialized financial tools, representing a significant leap in automated financial analysis Ding et al. (2024); Zhang et al. (2024). However, LLMs continue to encounter challenges in utilizing these tools’ outputs to generate high-quality analytical results. Two major challenges are the lack of specialized datasets for fine-tuning LLMs’ analytical capabilities, and the absence of systematic evaluation methods to quantify their performance in stock analysis. Furthermore, current research is constrained by LLMs’ reliance on historical data, such as GPT-4o’s dependence on its training data for generating responses Ni et al. (2024); Bhat and Jain (2024). This limitation in accessing real-time financial data and domain-specific information restricts them from fully capturing the dynamic nature of financial markets.
To address these limitations, we present three key contributions:
@@
  • Stocksis, a specialized dataset curated by industry experts to enhance LLMs’ stock analysis capabilities.
  • AnalyScore, a comprehensive evaluation framework to quantify improvements in analysis quality.
@@
  • FinSphere, an AI agent that can generate high-quality stock analysis reports in response to user queries.
Our experiments demonstrate that FinSphere, integrating real-time financial databases, specialized quantitative tools, and an instruction-tuned LLM, significantly outperforms both general and domain-specific LLMs, as well as existing agent-based systems. This superior performance persists even when baseline models are augmented with real-time background information and few-shot prompting, validating the effectiveness of our integrated approach to stock analysis.
2 Related Work
LLMs have emerged as powerful tools for stock analysis and trading Zhao et al. (2024); Li et al. (2023b). Research demonstrates their effectiveness in predicting stock prices and conducting market analysis Ni et al. (2024); Bhat and Jain (2024). Domain-specific models like InvestLM (Yang et al., 2023b) and GPT-InvestAR (Gupta, 2023) have been developed for investment analysis. Recent studies have also explored LLMs’ applications in financial anomaly detection Park (2024), portfolio evaluation Wu (2024), and financial statement analysis Kim et al. (2024), highlighting their potential to transform financial analysis.
Financial Datasets and Evaluation Metrics
Current financial datasets such as FinQA Chen et al. (2021), TAT-QA Zhu et al. (2021), and FLARE Xie et al. (2023) lack comprehensive stock analysis coverage, focusing instead on general financial tasks. Although BloombergGPT Wu et al. (2023) and FinRL Liu et al. (2021a) focus on trading strategies, and benchmarks like CFBenchmark Lei et al. (2023) cover broader financial tasks, a gap still exists in the availability of specialized stock analysis datasets. Additionally, financial text evaluation predominantly relies on general metrics such as BLEU Papineni et al. (2002) and ROUGE Rouge (2004), highlighting the necessity for domain-specific evaluation methods developed with financial expertise.
Instruction Tuning and Tool Integration
Financial LLMs have advanced through domain-specific instruction tuning, as demonstrated by InvestLM Yang et al. (2023b) and BloombergGPT Wu et al. (2023). Integration with specialized tools has also progressed, exemplified by FinGPT Yang et al. (2023a) with financial APIs, XBRL-Agent Han et al. (2024) with financial calculators, and the FinOps framework Li et al. (2023a). While these academic developments have improved financial applications Zhang et al. (2023b); Chen et al. (2023), they rely on historical data and fundamental tools. To address these limitations, we propose FinSphere Agent, leveraging our real-time database and advanced quantitative tools, as detailed in Section 4.
3 Stocksis and AnalyScore
Stock market analysis is becoming increasingly complex, necessitating the integration of diverse data sources and sophisticated analytical approaches. While LLMs demonstrate potential in transforming financial analysis, there are two critical gaps in the current landscape: the lack of high-quality training data for developing LLMs’ stock analysis capabilities, and the absence of standardized evaluation frameworks for assessing the quality of AI-generated stock analyses. This section introduces two significant contributions to address these gaps: Stocksis, a comprehensive dataset specifically designed to enhance LLMs’ stock analysis capabilities and AnalyScore, a systematic evaluation framework for assessing stock analysis reports.
@@ 3 Stocksis and AnalyScore
Table 1: An abbreviated example of Stocksis. The complete content is detailed in Table 6. Guide LLM to provide a comprehensive analysis based on specialized analyses returned from quantitative tools. The average cost of comprehensive analysis written by experts is \$10 per query, with detailed cost breakdowns available in Appendix G.
@@
3.1 Stocksis: A High-Quality Dataset for Enhancing LLMs’ Stock Analysis Capabilities
@@
Stocksis comprises 5,000 meticulously curated training pairs, with 50 pairs  [^1] The 50 pairs from Stocksis have been anonymized and are available at https://anonymous.4open.science/r/Stocksis-BD25/ available in the open-source release for research and development purposes. An abbreviated example is shown in Table 1 and the complete content of the same sample is detailed in Table 6. Each training sample consists of two key components:
@@
  • Prompt with Background Information (input): A complete analytical prompt that includes aggregated outputs from multiple quantitative analysis tools (averaging six distinct tools per sample) as background information. The background information covers volume-price analysis, technical indicators, and other market metrics. Each prompt is rigorously crafted to guide the model in performing specific analytical tasks while leveraging the provided background information. The average length is 4,000 words.
@@
  • Expert-Edited Analysis (label): In-depth analytical reports responding to the prompt’s requirements while effectively utilizing the background information, averaging 3,000 words per report. These expert-crafted analyses provide both overall market assessments and detailed rationales demonstrating how to effectively interpret various quantitative indicators.
@@
The dataset is meticulously compiled through a structured process leveraging our company’s expertise in stock market analysis. The quantitative tools used for data collection are well-established products from our company, which specializes in providing stock analysis recommendations to individual investors. The data collection process involved two key phases:
@@
    Prompt and Background Information Generation: Expert analysts carefully select appropriate analytical tools based on specific stock analysis queries, generating quantitative analyses as background information. These analysts subsequently craft targeted prompts that incorporate this background information to guide the analysis process.
@@
    Comprehensive Analysis Creation: A panel of 10 seasoned stock analysts develops detailed analysis reports in response to these prompts. Each analyst contributes their expertise to synthesize the background information into coherent insights that directly address the prompt requirements. This thorough process spans approximately three months, ensuring rigorous review and quality control of all analyses.
The release of Stocksis addresses a critical gap in AI financial analysis. While LLMs exhibit general capabilities, they struggle to integrate multiple quantitative signals into coherent market analysis. Current datasets primarily focus on either price data or news sentiment, lacking examples that combine structured prompts and expert reasoning. By publishing this dataset, we aim to facilitate the development of AI models capable of providing sophisticated, tool-based stock analysis, making quality financial insights more accessible.
AnalyScore is an innovative evaluation framework designed to assess the quality of stock analysis reports, developed by industry experts combining traditional stock analysis evaluation principles with LLM-related knowledge. This framework implements a two-tier evaluation system to ensure both fundamental quality standards and detailed analytical excellence.
Framework Structure
The evaluation process consists of two main components:
  1. 1.
    Priori Eligibility Check (shown in Table 4 in Appendix): A mandatory preliminary assessment comprising six essential criteria that must be met with 100% compliance prior to proceeding to detailed evaluation. These criteria ensure adherence to basic quality standards in (1) Conclusion Structure; (2) Logical Consistency; (3) Factual Support; (4) Data Timeliness; (5) Analytical Dimensions; (6) Neutral Language
  2. 2.
    Detailed Evaluation (shown in Table 5 in Appendix): A comprehensive scoring system across four key dimensions, totaling 100 points:
    • Conclusion (20 points): Evaluate the clarity and personalization of investment recommendations
    • Content (45 points): Assess professional analysis quality and logical consistency
    • Expression (15 points): Examine structural organization and language clarity
    • Data (20 points): Measure the breadth and depth of data utilization
This structured approach ensures a thorough and systematic evaluation of stock analysis reports, combining both qualitative standards and quantitative metrics to provide a comprehensive assessment of report quality. Currently, AnalyScore is employed solely by human experts in stock analysis, but we plan to design detailed prompts that enable LLMs to replace human experts in evaluating analysis reports using AnalyScore in the future.
@@ 4 FinSphere Agent
Figure 1: FinSphere Agent workflow and components interaction.
@@ 4.2 Instruction Tuning
We have conducted instruction fine-tuning on Qwen2-72B using our expert-curated Stocksis dataset to enhance the model’s financial analysis capabilities. The Stocksis dataset, as introduced in Section 3.1 comprises 5,000 high-quality training pairs, each consisting of structured prompts with comprehensive quantitative tool outputs and corresponding expert-authored analyses. The fine-tuning process employs the LoRA (Low-Rank Adaptation) Hu et al. (2021), which has enabled efficient parameter updates while maintaining the model’s general capabilities. Through this approach, we have trained the model to interpret outputs from multiple quantitative tools, integrate diverse analytical perspectives, and generate well-structured reports following professional analytical patterns.
@@ 4.3 Overall Workflow
FinSphere operates through a systematic, multi-stage process to generate comprehensive financial analyses. Upon receiving a user query, FinSphere first employs chain-of-thought (CoT) reasoning to decompose the analytical request into structured subtasks and identifies the appropriate quantitative tools required for each component.
Following task decomposition, the selected quantitative tools independently access our real-time financial database. Each tool retrieves the most current market data and information pertinent to its analytical domain, generating specialized analyses that reflect up-to-the-minute market conditions. These analyses range from technical indicators to fundamental metrics, providing a multi-dimensional view of the comprehensive analysis.
The final stage involves our Stocksis-tuned model, which serves as an expert analyst. The model receives all specialized analyses as input and synthesizes them into a cohesive, high-quality response. Through instruction fine-tuning on the Stocksis dataset, the model has developed sophisticated capabilities in interpreting quantitative outputs and generating professional-grade financial analyses. This integrated workflow ensures that FinSphere’s responses combine the precision of quantitative analysis with the nuanced understanding of expert financial reasoning, all while maintaining real-time relevance.
@@ 5 Evaluation
Qualification Conclusion Content Expression Data Total
Rate (Score: 20) (Score: 45) (Score: 15) (Score: 20) Score
GPT-4o 91% 9.85 26.12 12.44 18.20 66.61
GPT3.5 78% 7.95 21.05 10.15 14.30 53.45
Qwen2-72B 81% 8.15 22.55 10.55 14.95 56.20
InvestLM 75% 8.40 23.10 11.25 15.75 58.50
FinGPT 62% 6.80 18.55 8.95 10.75 40.05
FinRobot 81% 9.10 24.05 11.55 16.35 61.05
FinMem 84% 9.90 25.95 12.85 18.85 67.55
FinSphere 89% 9.95 27.16 14.87 18.90 70.88
Table 2: Human experts use AnalyScore to evaluate 100 responses generated by 8 models. The scores shown are averages across 100 evaluations. The average cost for expert evaluation is \$10 per response, with detailed cost breakdowns available in Appendix G.
Given FinSphere’s integration with real-time financial databases and proprietary quantitative tools, it possesses analytical capabilities that extend beyond those of general-purpose LLMs. Traditional performance comparisons between FinSphere and general LLMs present inherent challenges, primarily due to the latter’s inability to access real-time financial data and domain-specific information. For example, GPT-4o typically acknowledges its limitations with responses like "As an AI language model with knowledge cut-off in October 2023, I don’t have access to real-time stock information." To demonstrate FinSphere’s enhanced capabilities while ensuring a fair comparison, we have implemented a comprehensive experimental design.
@@
Our comparative analysis evaluates three categories of models: single LLM, Agent-based systems, and FinSphere. For LLM-based models, we test proprietary models (GPT-4o, GPT3.5), open-source model (Qwen2-72B), and domain-specific models (InvestLM Yang et al. (2023b), FinGPT Yang et al. (2023a)), all using chain-of-thought prompting with few-shot examples and relevant background information (detailed in Appendix D). For Agent-based systems, we evaluate FinMem Yu et al. (2024) and FinRobot Yang et al. (2024), employing simplified prompts with few-shot examples and background information, like the input of Stocksis. Finally, we evaluate FinSphere through direct user queries, leveraging its integrated real-time database and quantitative tools (detailed in Appendix I).
5.1 Generation Contrast
Comprehensive results in Table 2 evaluated by experts using the AnalyScore framework demonstrate FinSphere’s superior performance across all assessment dimensions, achieving an overall score of 70.88 out of 100. This performance notably surpasses both traditional LLM-based approaches and other agent-based systems, with FinMem and GPT-4o following at 67.55 and 66.61 respectively. The evaluation reveals a clear performance hierarchy. Generally, specialized agent-based systems consistently outperform standalone language models (except GPT-4o), while general-purpose LLMs show moderate performance and domain-specific LLMs like FinGPT (40.05) demonstrate relatively limited capabilities. The results validate the effectiveness of FinSphere’s integrated approach, which combines real-time data access, quantitative tools and the Stocksis-tuned LLM.
It is noteworthy that general LLMs require extensive exemplars in their prompts to generate appropriate stock analysis outputs. This requirement significantly increases the input token count, resulting in higher operational costs for GPT-4o and restrictions on small context-window LLM. In contrast, FinSphere’s architecture and instruction fine-tuning obviate the need for such verbose prompts, allowing it to achieve superior performance with substantially fewer input tokens. Furthermore, FinSphere is scheduled for public release with free access in December 2024, with detailed release information available in Appendix H.

### v2

# FinSphere, a Real-Time Stock Analysis Agent Powered by Instruction-Tuned LLMs and Domain Tools

@@ Abstract
Current financial large language models (FinLLMs) struggle with two critical limitations: the absence of objective evaluation metrics to assess the quality of stock analysis reports and a lack of depth in stock analysis, which impedes their ability to generate professional-grade insights. To address these challenges, this paper introduces FinSphere, a stock analysis agent, along with three major contributions: (1) AnalyScore, a systematic evaluation framework for assessing stock analysis quality, (2) Stocksis, a dataset curated by industry experts to enhance LLMs’ stock analysis capabilities, and (3) FinSphere, an AI agent that can generate high-quality stock analysis reports in response to user queries. Experiments demonstrate that FinSphere achieves superior performance compared to both general and domain-specific LLMs, as well as existing agent-based systems, even when they are enhanced with real-time data access and few-shot guidance. The integrated framework, which combines real-time data feeds, quantitative tools, and an instruction-tuned LLM, yields substantial improvements in both analytical quality and practical applicability for real-world stock analysis.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in comprehending and processing natural language, extending their influence across various domains, including finance (Li et al., 2023b). By leveraging their language comprehension capabilities, these models have exhibited exceptional performance in various financial applications, including sentiment analysis (Liu et al., 2024b; Zhang et al., 2023a) and information extraction from unstructured financial texts (Li et al., 2024; Huang et al., 2023). The advent of finance-specific LLMs such as FinBERT (Yang et al., 2020; Liu et al., 2021), BloombergGPT (Wu et al., 2023), and PIXIU (Xie et al., 2023) has further enhanced the capacity to process financial data effectively. These advancements have laid the foundation for developing more sophisticated financial analysis tools and shifted how investors interact with market data (Krause, 2023; Nie et al., 2024). These AI-powered systems have broadened access to professional financial insights, allowing retail investors to benefit from advanced analysis once reserved for institutions.
@@
As LLM technology continues to evolve, there is a growing expectation for these models to handle more complex financial tasks, particularly in stock analysis and real-time financial question-answering (Yang et al., 2023b; Zhao et al., 2024). This has led to the development of tool-augmented agents that integrate LLMs’ natural language understanding with specialized financial tools, significantly enhancing automated financial analysis and interactive question-answering capabilities (Ding et al., 2024; Zhang et al., 2024). However, LLM-based financial QA systems still face substantial challenges in effectively interpreting and utilizing the outputs of these tools to generate high-quality analytical responses. Two primary obstacles include the absence of systematic evaluation frameworks to assess their performance in stock analysis, as well as the lack of specialized datasets for fine-tuning LLMs’ analytical reasoning capabilities. Moreover, existing research is constrained by LLMs’ heavy reliance on historical data, such as GPT-4o’s dependence on its pre-trained knowledge for generating responses (Ni et al., 2024; Bhat and Jain, 2024). This limitation in accessing and processing real-time financial data and domain-specific information restricts their ability to fully capture the dynamic and evolving nature of financial markets, posing a critical challenge for real-time financial question-answering systems.
We address these limitations through three contributions:
  • AnalyScore: A comprehensive evaluation framework designed to systematically assess the accuracy, relevance, and analytical depth of LLM-driven stock question-answering.
@@
  • Stocksis: A specialized dataset curated by industry experts to enhance LLMs’ financial question-answering and stock analysis capabilities.
@@
  • FinSphere: A real-time stock question-answering agent capable of generating high-quality stock analysis reports in response to user queries.
@@
Table 1: An abbreviated example of Stocksis. The complete content is detailed in Table 5. Guide LLM to provide a comprehensive analysis based on specialized analyses returned from quantitative tools. The average cost of comprehensive analysis written by experts is \$10 per query, with detailed cost breakdowns available in Appendix F.
Our experiments demonstrate that FinSphere, by integrating real-time financial databases, specialized quantitative tools, and an instruction-tuned LLM optimized for financial question-answering, significantly outperforms both general-purpose and domain-specific LLMs, as well as existing agent-based systems. This superior performance holds even when baseline models are augmented with real-time background information and few-shot prompting, validating the effectiveness of our integrated approach to real-time financial question-answering and stock market analysis.
2 AnalyScore and Stocksis
Stock market analysis is becoming increasingly complex, necessitating the integration of diverse data sources and sophisticated analytical approaches. While LLMs show promise in financial analysis, two key gaps remain: the lack of standardized evaluation frameworks for AI-generated stock analyses and the scarcity of high-quality training data. To address these, we introduce AnalyScore, a systematic evaluation framework, and Stocksis, a comprehensive dataset designed to enhance LLMs’ stock analysis capabilities.
To address the lack of systematic evaluation standards in stock analysis, we introduce AnalyScore, a domain-specific framework co-developed with industry experts. It combines established financial assessment criteria with insights from LLM-generated content evaluation.
AnalyScore employs a structured scoring system across four dimensions, with a total score of 100 points:
  • Conclusion (20 pts): Clarity, relevance, and personalization of investment recommendations.
  • Content (45 pts): Depth, coherence, and professionalism of analytical reasoning.
  • Expression (15 pts): Organization, fluency, and linguistic clarity.
  • Data (20 pts): Breadth, accuracy, and effective use of supporting quantitative data.
This multi-dimensional evaluation ensures comprehensive coverage of both qualitative judgment and quantitative rigor. At present, AnalyScore is used exclusively by human experts to assess model-generated analyses. In future work, we aim to integrate AnalyScore into automatic evaluation pipelines by prompting LLMs to emulate expert scoring behaviors.
@@ 2.1 AnalyScore: A Comprehensive Evaluation Framework for Stock Analysis Reports
2.2 Stocksis: A High-Quality Dataset for Enhancing LLMs’ Stock Analysis
To evaluate LLMs’ capabilities in stock analysis, we first assess GPT-4o’s responses using AnalyScore (Section 4.1), providing it with extensive background knowledge including market data and quantitative indicators. Despite these supports, its outputs often exhibit reasoning inconsistencies, shallow financial insights, and occasional misinterpretations of market trends—revealing the challenges LLMs face in synthesizing complex financial information into coherent, actionable analysis. To address these limitations, we collaborate with industry experts to iteratively refine GPT-4o’s outputs, correcting errors and enriching them with deeper reasoning. This expert-guided process results in Stocksis, a high-quality dataset that bridges automated generation and professional-grade analysis, offering structured supervision to improve LLM performance in real-world financial contexts.
@@ 2.2 Stocksis: A High-Quality Dataset for Enhancing LLMs’ Stock Analysis
Stocksis comprises 5,000 meticulously curated training pairs, with part of them [^1] Open-sourced Stocksis has been anonymized and is available at https://anonymous.4open.science/r/Stocksis-BD25/ available in the open-source release for research and development purposes. An abbreviated example is shown in Table 1, and the complete content of the same sample is detailed in Table 5. Each training sample consists of two key components:
@@
  • Prompt with Background Information (input): A complete analytical prompt that includes aggregated outputs from multiple quantitative analysis tools (averaging six tools per sample) as background information. The background information covers volume-price analysis, technical indicators, and other market metrics. Each prompt is rigorously crafted to guide the model in performing analytical tasks while leveraging the provided background information. The average length is 4,000 words.
@@
  • Expert-Edited Analysis (label): In-depth analytical reports responding to the prompt’s requirements while effectively utilizing the background information, averaging 3,000 words per report. Due to the particularity of the stock analysis task, there is no standard answer to this task. Therefore, our industry experts provide a high-quality reference analysis result for this task by evaluating the overall market, providing detailed reasons, and demonstrating how to effectively interpret various quantitative indicators.
@@
Stocksis is constructed through a structured pipeline grounded in our company’s expertise in retail-oriented stock analysis. Data collection involves two main stages:
@@
    Prompt and Background Generation: Expert analysts select suitable quantitative tools for specific stock queries and craft prompts enriched with structured analytical outputs.
@@
    Analysis Refinement: GPT-4o generates initial responses, which are then iteratively reviewed and improved by a panel of 10 senior analysts to ensure accuracy, coherence, and domain relevance. This process spans over three months to ensure high-quality outputs.
Stocksis addresses a key gap in financial NLP: the lack of datasets combining structured prompts with expert-refined reasoning. Unlike existing resources focused on price or sentiment data, it enables the fine-tuning of general-purpose LLMs to enhance their capability in structured, high-quality financial analysis for real-world applications.
@@ 3 FinSphere Agent
Figure 1: This diagram illustrates the overall workflow of the FinSphere Agent, detailing how different components interact to facilitate real-time stock analysis.
Conclusion Content Expression Data Total
(Score: 20) (Score: 45) (Score: 15) (Score: 20) Score
GPT-4o 9.85 26.12 12.44 18.20 66.61
Deepseek-v3 9.52 25.30 12.75 16.85 64.42
GPT3.5 7.95 21.05 10.15 14.30 53.45
Qwen2-72B 8.15 22.55 10.55 14.95 56.20
InvestLM 8.40 23.10 11.25 15.75 58.50
FinGPT 6.80 18.55 8.95 10.75 40.05
FinRobot 9.10 24.05 11.55 16.35 61.05
FinMem 9.90 25.95 12.85 18.85 67.55
FinSphere 9.95 27.16 14.87 18.90 70.88
Table 2: Human experts use AnalyScore to evaluate 100 responses generated by 8 models. The scores shown are averages across 100 evaluations. The average cost for expert evaluation is \$10 per response, with detailed cost breakdowns available in Appendix F.
@@ 3.2 Instruction Tuning
We adopt full-parameter instruction tuning to adapt Qwen2-72B into a domain-specialized financial analysis agent. Using our expert-curated Stocksis dataset—comprising 5,000 structured training pairs of quantitative tool outputs and corresponding expert-written analyses (Section 2.2)—we guide the model to learn not just domain knowledge, but also the task format, reasoning flow, and response style required for professional stock reporting. Unlike parameter-efficient methods, full instruction tuning enables the model to internalize complex analytical patterns and generate coherent, multi-dimensional reports.
[…]
@@ 3.3 Overall Workflow
FinSphere follows a structured, multi-stage workflow to generate real-time financial analyses. Upon receiving a user query, it first applies chain-of-thought (CoT) reasoning to decompose the task into interpretable subtasks and determine which domain-specific quantitative tools are needed. These tools then independently access a real-time financial database to retrieve the latest market data, including technical indicators, capital flows, and fundamental metrics, each generating specialized outputs tailored to their analytical focus.
In the final stage, FinSphere’s instruction-tuned LLM—trained on the Stocksis dataset—serves as an expert analyst. It synthesizes the multi-source outputs into a coherent, structured report aligned with professional financial standards. This integrated architecture enables FinSphere to combine the precision of automated quantitative analysis with the contextual depth of expert reasoning, ensuring responses are both analytically sound and up-to-date.
@@ 4 Evaluation
Given FinSphere’s integration with real-time financial databases and proprietary quantitative tools, it possesses analytical capabilities that extend beyond those of general-purpose LLMs. Performance comparisons between FinSphere and general LLMs present inherent challenges, primarily due to the latter’s inability to access real-time financial data and domain-specific information. For example, GPT-4o typically acknowledges its limitations with responses like ”As an AI language model with knowledge cut-off in June 2024, I don’t have access to real-time stock information.” To demonstrate FinSphere’s enhanced capabilities while ensuring a fair comparison, we have implemented a comprehensive experimental design.
Group 1 & 2 Group 1 & 3 Group 1 & 4 Group 2 & 3 Group 2 & 4 Group 3 & 4 Average
Conclusion 79.36 73.90 79.51 81.59 85.30 89.63 81.55
Content 93.77 71.45 94.25 77.61 73.49 74.99 80.93
Expression 88.30 91.65 90.81 83.12 77.30 82.86 85.67
Data 84.97 85.03 75.31 80.80 79.16 84.81 81.68
Total 78.90 87.70 74.55 77.28 81.40 73.16 78.83
Table 3: Average Kendall’s Tau across 100 queries for different annotator groups rating. Reported as %
@@
We evaluate three categories of models: (1) single LLMs, including proprietary (GPT-4o, GPT-3.5), open-source (Qwen2-72B, Deepseek-v3 (Liu et al., 2024a)), and domain-specific models (InvestLM (Yang et al., 2023b), FinGPT (Yang et al., 2023a)); (2) agent-based systems, including FinMem (Yu et al., 2024) and FinRobot (Yang et al., 2024); and (3) our proposed FinSphere. All LLMs use chain-of-thought prompting with few-shot examples and relevant background information (Appendix C), while agents receive simplified prompts similar to Stocksis’s inputs. FinSphere is evaluated directly via user queries, utilizing its integrated tools (Appendix H). For all models, we set a maximum output length of 8K tokens and a temperature of 0.5.
4.1 Performance Analysis
The expert evaluation results presented in Table 2 demonstrate FinSphere’s superior performance across all assessment dimensions, achieving an overall score of 70.88 out of 100. This surpasses both traditional LLM-based approaches and other agent-based systems, with FinMem and GPT-4o following at 67.55 and 66.61, respectively. The evaluation reveals a clear performance hierarchy: agent-based systems generally outperform standalone language models (except GPT-4o), while general-purpose LLMs show moderate performance and domain-specific LLMs such as FinGPT (40.05) demonstrate relatively limited capabilities. These results validate the effectiveness of FinSphere’s integrated approach, which combines real-time data access, quantitative tools, and a Stocksis-tuned LLM, enabling more precise and insightful stock analysis. More analysis is shown in Appendix B.
4.2 Evaluation Consistency
We assess the reliability of human evaluation using Kendall’s Tau rank correlation to measure inter-group agreement. Forty industry experts are divided into four groups of ten, each producing consensus scores for all model responses. Pairwise Kendall’s Tau is computed across groups based on rankings of 100 queries (Table 3).
Results show strong agreement, with most correlations exceeding 80% and values ranging from 71.45 to 94.25. While Content shows the most variation, Expression and Data maintain stable consistency. These findings indicate that, despite inherent subjectivity, our evaluation process yields robust and reliable assessments of model performance.